# A comparative analysis of metamodels for lumped cardiovascular models, and pipeline for sensitivity analysis, parameter estimation, and uncertainty quantification

### v1

@@ 3.1.3 Gaussian processes
Gaussian Processes (GPs) are a powerful statistical tool for modeling and understanding complex datasets [56]. At their core, GPs are a type of Bayesian inference method that provides a flexible approach to regression and classification problems. They excel in situations where the underlying structure of the data is unknown or difficult to specify with traditional parametric models.
@@
In this work, we employed the Python library scikit-learn [57] to implement our Gaussian Process model.
@@ 3.2 Data preparation and hyperparameter selection
In this work, synthetic data from each model are the basis to train the metamodels. The generated data are split into training and testing sets. For all models, the data are generated by implementing the sampling method described by Saltelli’s algorithm [58], which is a quasi-Monte Carlo approach, mainly used to calculate the Sobol sensitivity indices [59]. The open-source Python library SALib [60] was implemented. The maximum computational cost for this approach is $N_{s}=(2d+2)N$, where $N_{s}$ represents the number of simulations or model evaluations, $d$ signifies the model’s input variables, and $N$ denotes the sample size.
@@
For the polynomial chaos expansion, an adaptive algorithm iteratively removes polynomial terms $\alpha_{j}$ with a low relative contribution, with $\mathcal{A}_{rem}$ the subset of polynomial terms to be removed,
$\mathcal{A}_{rem}=\{\|\alpha_{j}\|\geq\epsilon\max_{\alpha\in\mathcal{A}_{tot}}\boldsymbol{(\alpha)}\}$ (7)
with $\mathcal{A}_{tot}$ the total set of polynomials, with the error term $\epsilon$ set to $10^{-4}$.
@@ 4 Metamodel building and assessment
For all the models, the performances of neural networks, PCE, and Gaussian processes are compared. In all cases, 20% of the data is kept for testing, and the $Q^{2}$ metric (Eq. 8) is measured using different training set sizes along with the maximum error. This is done to see the effect of the training data size on the metamodels’ performance.
@@
where $\mathbf{\bar{Y}}$ is the mean of the true values $\mathbf{Y}_{i}$.
@@ 4.1 Model 1
Table 2 presents the different types of model inputs as well as model outputs, while Q2 for all outputs is plotted in figure 5.
@@
Other Organ Resistance ($R_{OO}$) Mean Aortic Pressure ($MAP$)
Digestive Organ Resistance ($R_{DO}$) Cardiac Output ($CO$)
Portal Vein Resistance ($R_{PV}$) Portal Vein Pressure ($P_{PV}$)
Hepatic Artery Resistance ($R_{HA}$) Vena Cava Pressure ($P_{VC}$)
Active Left Ventricular Elastance ($Ea_{LV}$) Portal Vein Flow ($Q_{PV}$)
Baseline Left Ventricular Elastance ($Eb_{LV}$) Hepatic Artery Flow ($Q_{HA}$)
Post-procedural Injected Volume ($V_{inj}$)
Percentage of removed liver tissue ($Hpx$)
Post-procedural Heart Rate Change ($\Delta HR$)
Figure 5: Q2 metric using the testing dataset for the first model.
@@ 4.2 Model 2
For this model, we have selected the ten most influential and clinically relevant inputs to the outputs based on the Morris Sensitivity Method [61]. The Morris method performs a qualitative sensitivity analysis primarily suited for models with numerous independent input parameters. Given that the current model initially consists of over 40 independent inputs, many of which have minimal impact on the outcomes, the Morris method offers a cost-effective means to qualitatively rank them, in contrast to the Sobol indices method. With regard to the model outputs, the selection is based on the expertise of clinicians. The detailed selection of the model inputs and outputs for the metamodels’ structure is presented in Table 3
@@
Figure 7: Q2 metric using the testing dataset for the second model.
@@ 4.3 Model 3
Since the output is a time-dependent variable, a different metamodeling technique need to be used. Fourier transform was initially our first candidate, however, the signals are not periodic and most them are early cut and most of the frequencies knowledge is lost.
@@
Therefore, specifically for this model, we only implemented LSTM model in Tensorflow. Figure 9 shows samples from the test sets and their corresponding LSTM predictions, which clearly shows good agreement. The two cases are chosen due to their different signal behaviour corresponding to the diseased level of the patient.
Figure 9: Two testing samples of RNN predictions of the contrast agent concentration. On the left, one can find a signal representing a healthy person, while on the left a diseased person.
@@ 5.1 Sensitivity analysis
Variance-based sensitivity analysis (Sobol indices), which is a global sensitivity analysis technique, is chosen in this work. When PCE is chosen as the surrogate model, Sobol indices can be derived directly from the coefficients of the polynomial terms [17], whereas for NNs and GPs, a classical Monte Carlo approach is used [59]. In this section, we compare the total Sobol indices for the PAH model, first calculated for its 0D version using Saltelli’s method, and then for its surrogate model built with neural networks.
@@
Figure 10: Sobol indices for the second cardiovascular model. On the left, one can find the total Sobol indices using the 0D model, and on the left using neural network predictions.
@@ 5.2 Parameter estimation
where $\mathbf{X}=\boldsymbol{\sigma(\psi)}$, $\boldsymbol{\sigma}$ is a function designed to constrain to input space and is chosen as a variant of the sigmoid function, and $\boldsymbol{\psi}$ is a dummy variable. This construction is chosen to ensure that the estimated parameters falls within physiological ranges. n problem, $\boldsymbol{\psi}$ is initialized randomly and then updated through the minimization process with Adam optimization algorithm. An example of a solved inverse problem is shown in figure 11 where 50000 Adam iterations are performed taking nearly 2 minutes. All parameter estimations converged.
@@
where $X_{est}$ si the value of the estimated parameter, $X_{ref}$ the reference value, and $X_{max},\ X_{low}$ are the physiological limits of the parameter.
@@ 5.3 Uncertainty quantification
Due to the relatively low computational cost of the NN emulator to solve inverse problems and forward predictions, patient specific UQ was performed using a two-stepped, Monte Carlo type approach. First, an inverse UQ is performed, whereby patient data and their assumed uncertainty (Table 6) is sampled, and the inverse problem (as described in 5.2) is solved for each individual sample. Note that the inverse problem is solved with a NN that is trained on the pre-intervention model, i.e. without the Potts shunt in place. Convergence to a stable distribution of optimized parameters is monitored by plotting the mean and variance over the number of samples solved (Fig 12). Once convergence is achieved, the resulting optimized parameter distributions are then used to predict the hemodynamic outcome of the placement of a Potts shunt.
Table 6: Measured data for an example patient, as published in earlier work [6]. All measurement errors are assumed independent and uniformly distributed.
@@
$P_{ao,dias}$ 53 mmHg 2.5% [62]
$P_{ao,mean}$ 69 mmHg 2.5 % [62]
$P_{ao,sys}$ 94 mmHg 2.5 % [62]
$P_{pa,dias}$ 67 mmHg 2.5 % [62]
$P_{pa,mean}$ 85 mmHg 2.5 % [62]
$P_{pa,sys}$ 112 mmHg 2.5 % [62]
$SV_{LV}$ 51 mL 8 % [63]
$SV_{RV}$ 51 mL 8 % [63]
[…]
@@ 6.1 Computational time
The computational time of the different metamodels employed is discussed in this section as shown in Table 7. For the building of metamodels, it was observed that neural networks have the fastest training time in the case of a big dataset ($>30K$ data points). The computational time was always in the order of few minutes when training with the full dataset. However, this cost goes to nearly an hour for PCE and around 12 hours for GPs, when running on the same laptop. This is the price paid by GPs in order to offer a natural way to quantify uncertainty in predictions, providing the predicted mean but also a confidence interval, representing the uncertainty at each prediction point. This is particularly valuable in fields like Bayesian optimization, where understanding uncertainty is crucial. It should be noted that when training on a small part of the dataset, 3K points for example, GPs and PCE have training times in the order of minutes.
Once an emulator is trained, obtaining Sobol indices using the Saltelli method [58] is a relatively fast process for NN and GPs (few minutes), due to the low computational cost per emulator evaluation. PCE has the advantage that Sobol indices are readily available when the model has been trained, with no additional computational cost.
@@
mInverse problem $<$ A minute Requires additional setup Requires additional setup
@@ 6.2 Ensuring physiological validity in metamodels
For prediction purposes, the challenge is to ensure that the outputs of the metamodel adhere to physiological constraints. This can be approached in two main ways. One method involves training the metamodel exclusively on data that is known to fall within physiological ranges. By doing so, the model learns to generate predictions that are inherently physiologically plausible. However, this approach limits the model’s applicability to the specific ranges of inputs and outputs included in the training data. The second method involves using post-prediction filtering. In this approach, the metamodel generates predictions based on a broader set of data, but these predictions are subsequently checked and corrected to ensure they fall within physiological limits. While this method helps ensure that final outputs are valid, it does not prevent the generation of non-physiological outputs during the modeling process itself.
In the context of SA, which aims to understand how variations in input parameters affect output variability, the approach can differ based on the tools used. One effective strategy proposed in [18] is to build a metamodel based on PCE using a training set restricted to input-output pairs that are physiologically valid. This approach ensures that the PCE provides Sobol indices naturally, reflecting the sensitivity of outputs to inputs while avoiding biases from non-physiological data. For those who may not have access to PCE or are unfamiliar with it, another approach is to create a general metamodel, such as one based on GP or NN, using a dataset that includes both physiological and non-physiological pairs. Sobol indices can then be computed using the classical Monte-Carlo approach as shown in Sec. 5.1.
@@ 6.3 Parameter estimation and inverse problems
While outside of the scope of this work, in some cases there are several values or combinations of inputs that lead to the same output, i.e., the chosen set of input parameters is unidentifiable. Depending on the research question, there are several approaches to deal with unidentifiability. First is a Monte Carlo type approach, where the inverse problem is solved several times with different random initializations, leading to a distribution of input parameters [64]. Some methods leverage the sensitivity indices to derive an orthogonal set of input parameters [65, 66], while other methods allow for a parametric description of non-identifiable parameter combinations [67]. Alternatively, a formal identifiability analysis is performed, iteratively reducing the input space by fixing unidentifiable parameters to produce a reduced, identifiable set; such as profile likelihood analysis [68, 69], diaphony [70, 71] or determining intra-class correlation [72]. However, these methods can be prohibitive due to their computational cost, which can be overcome by using surrogate models.
@@ 6.4 Conclusion
To conclude, overall NN present advantages in terms of implementation and run-time cost-effectiveness to perform the whole pipeline of studying and performing patient-specific modelling.

### v2

@@ 3.1.3 Gaussian processes
Gaussian Processes (GPs) are a powerful statistical tool for modeling and understanding complex datasets [marrel2024probabilistic]. At their core, GPs are a type of Bayesian inference method that provides a flexible approach to regression and classification problems. They excel in situations where the underlying structure of the data is unknown or difficult to specify with traditional parametric models.
@@
In this work, we employed the Python library scikit-learn [scikit-learn] to implement our Gaussian Process model.
@@ 3.2 Data preparation and hyperparameter selection
In this work, synthetic data from each model are the basis to train the metamodels. The generated data are split into training and testing sets. For all models, the data are generated by implementing the sampling method described by Saltelli’s algorithm [saltelli2004sensitivity], which is a quasi-Monte Carlo approach, mainly used to calculate the Sobol sensitivity indices [sobol2001global]. The open-source Python library SALib [herman2017salib] was implemented. The maximum computational cost for this approach is $N_{s}=(2d+2)N$, where $N_{s}$ represents the number of simulations or model evaluations, $d$ signifies the model’s input variables, and $N$ denotes the sample size.
@@
For the polynomial chaos expansion, an adaptive algorithm iteratively removes polynomial terms $\alpha_{j}\Psi_{j}$ with a low relative contribution, i.e. where the coefficient $\alpha_{j}$ is below a certain threshold
$\mathcal{A}_{rem}=\{\|\alpha_{j}\|\leq\epsilon\max_{\alpha\in\mathcal{A}_{tot}}\boldsymbol{(\alpha)}\},$ (7)
with $\mathcal{A}_{tot}$ the total set of polynomials and $\mathcal{A}_{rem}$ the set of polynomials to be removed. The error term $\epsilon$ set to $10^{-4}$.
@@ 4 Metamodel building and assessment
The same training and testing datasets are used for neural networks, PCE, and GP, to ensure fair comparison. In all cases, 20% of the data is kept for testing, and the $Q^{2}$ metric (Eq. 8) is measured using different training set sizes (same datasets for all methods) along with the maximum error. This is done to see the effect of the training data size on the metamodels’ performance.
@@
where $\mathbf{\bar{Y}}$ is the mean of the true values $\mathbf{Y}_{i}$, and $n$ is the testing dataset size.
@@ 4.1 Model 1
Table 2 presents the different types of model inputs as well as model outputs. Note that some parameters which are commonly considered outputs (e.g., Cardiac Output, Mean Aortic Pressure) are used as input to the model. This is due to the workflow whereby several resistances are directly derived from measured data, as explained in [golse2021predicting]. The $Q^{2}$-metric for all outputs is plotted in figure 5.
@@
Active Right Atrial Elastance ($Ea_{RA}$) Mean Aortic Pressure, achieved ($MAP$)
Baseline Right Atrial Elastance ($Eb_{RA}$) Portal Vein Pressure, achieved ($P_{PV}$)
Active Right Ventricular Elastance ($Ea_{RV}$) Vena Cava Pressure, achieved ($P_{VC}$)
Baseline Right Atrial Elastance ($Eb_{RV}$) Cardiac Output, achieved ($CO$)
Active Left Atrial Elastance ($Ea_{LA}$) Hepatic Artery Flow ($Q_{HA}$)
Baseline Left Atrial Elastance ($Eb_{LA}$) Portal Vein Flow ($Q_{PV}$)
Active Left Ventricular Elastance ($Ea_{LV}$) Left Ventricular Ejection Fraction ($EF$)
Baseline Left Ventricular Elastance ($Eb_{LV}$) Mean Left Atrial Pressure ($MLAP$)
Vena Cava Compliance ($C_{vc}$) Peak Pulmonary Artery Pressure ($PPAP$)
Mean Aortic Pressure, target ($MAP_{m}$) Right Atrial End Diastolic Volume ($RAEDV$)
Portal Vein Pressure, target ($P_{PV,m}$) Right Ventricular End Diastolic Volume ($RVEDV$)
Vena Cava Pressure, target ($P_{VC,m}$) Left Atrial End Diastolic Volume ($LAEDV$)
Cardiac Output, target ($COm$)
Hepatic Artery Flow, target ($Q_{HA,m}$)
Portal Vein Flow, target ($Q_{PV,m}$)
Left Liver Volume ($M_{L}$)
Right Liver Volume ($M_{R}$)
Heart Rate ($HR$)
Figure 5: $Q^{2}$ metric using the testing dataset for the first model.
@@ 4.2 Model 2
For this model, we have selected the ten most influential and clinically relevant inputs to the outputs based on the Morris Sensitivity Method [morris1991factorial]. The Morris method performs a qualitative sensitivity analysis primarily suited for models with numerous independent input parameters. Given that the current model initially consists of over 40 independent inputs, many of which have minimal impact on the outcomes, the Morris method offers a cost-effective means to qualitatively rank them, in contrast to the Sobol indices method. With regard to the model outputs, the selection is based on the expertise of clinicians. The detailed selection of the model inputs and outputs for the metamodels’ structure is presented in Table 3
@@
Figure 7: $Q^{2}$ metric using the testing dataset for the second model.
@@ 4.3 Model 3
Since the output is a time-dependent variable, a different metamodeling technique needs to be used. Fourier transform was initially our first candidate, however, the signals are not periodic and most them are early cut and most of the frequencies knowledge is lost.
@@
Therefore, specifically for this model, we implemented an LSTM architecture in Tensorflow. The LSTM architecture consists in a few-to-sequence model outputting a sequence of size 200. In the architecture, only one LSTM block was implemented since it was enough to produce accurate predictions.
Figure 9 shows samples from the test sets and their corresponding LSTM predictions, which clearly shows good agreement. The two cases are chosen due to their different signal behavior corresponding to the diseased level of the patient.
Figure 9: Two testing samples of LSTM predictions of the contrast agent concentration. On the left, one can find a signal representing a healthy case (low vascular anomaly intensity $s_{0}$), while on the left a diseased case (higher vascular anomaly intensity $s_{0}$).
@@ 5.1 Sensitivity analysis
Variance-based sensitivity analysis (Sobol indices), which is a global sensitivity analysis technique, is chosen in this work. When PCE is chosen as the surrogate model, Sobol indices can be derived directly from the coefficients of the polynomial terms [sudret2008global], whereas for NNs and GPs, a classical Monte Carlo approach is used [sobol2001global]. In this section, we compare the total Sobol indices for the PAH model, first calculated for its 0D version using Saltelli’s method, and then for its surrogate model built with neural networks.
@@
Figure 10: Sobol indices for the second cardiovascular model. On the left, one can find the total Sobol indices using the 0D model, and on the middle using neural network predictions, and the absolute error is plotted on the right.
@@ 5.2 Parameter estimation
where $\mathbf{X}=\boldsymbol{\sigma(\psi)}$, $\boldsymbol{\sigma}$ is a function designed to constrain to input space and is chosen as a variant of the sigmoid function $\mathbf{S}$, and $\boldsymbol{\psi}$ is a dummy variable. This construction is chosen to ensure that the estimated parameters falls within physiological ranges. The function $\boldsymbol{\sigma(\psi)}$ can be written as:
$\boldsymbol{\sigma(\psi)}=\mathbf{X_{1}+(X_{2}-X_{1})S(\psi)},$ (11)
where $\mathbf{X_{1}}$ and $\mathbf{X_{2}}$ are the lower and upper bounds of the function. It should be noted that this addition does not ensure the validity of the solution of the inverse problem since some of the input parameters might not be uniquely identifiable. However, if there is a single local minimum in that specified range, it helps reaching that specific minimum instead of falling in a non physiological minimum.
To solve this minimization problem, $\boldsymbol{\psi}$ is initialized randomly and then updated through the minimization process with Adam optimization algorithm. An example of a solved inverse problem is shown in figure 11 where 50000 Adam iterations are performed taking nearly 2 minutes. All parameter estimations converged.
@@
where $X_{est}$ is the value of the estimated parameter, $X_{ref}$ the reference value, and $X_{max},\ X_{low}$ are the physiological limits of the parameter.
@@ 5.3 Uncertainty quantification
Due to the relatively low computational cost of the NN emulator to solve inverse problems and forward predictions, patient specific UQ was performed using a two-stepped, Monte Carlo type approach. First, an inverse UQ is performed, whereby patient data and their assumed uncertainty (Table 6) is sampled, and the inverse problem (as described in 5.2) is solved for each individual sample. Note that the inverse problem, as visualized in Figure 1, is solved with a NN that is trained on the pre-intervention model, i.e. without the Potts shunt in place, since the model should be optimized to pre-intervention data to reflect the patient state before treatment. Convergence to a stable distribution of optimized parameters is monitored by plotting the mean and variance over the number of samples solved (Fig 12). Once convergence is achieved, the resulting optimized parameter distributions are then augmented with the model parameters which describe the placement of the shunt, and are used to predict the hemodynamic outcome of the placement of a Potts shunt. A similar result could be achieved by sampling a wider input range for the parameters describing the shunt resistance, which would lead to neglible flow across the shunt. However, this would likely significantly increase the number of samples required to train the NN.
Table 6: Measured data for an example patient, as published in earlier work [pant2022multiscale]. All measurement errors are assumed independent and uniformly distributed.
@@
$P_{ao,dias}$ 53 mmHg 2.5% [edwardsmanual]
$P_{ao,mean}$ 69 mmHg 2.5 % [edwardsmanual]
$P_{ao,sys}$ 94 mmHg 2.5 % [edwardsmanual]
$P_{pa,dias}$ 67 mmHg 2.5 % [edwardsmanual]
$P_{pa,mean}$ 85 mmHg 2.5 % [edwardsmanual]
$P_{pa,sys}$ 112 mmHg 2.5 % [edwardsmanual]
$SV_{LV}$ 51 mL 8 % [mercado2017transthoracic]
$SV_{RV}$ 51 mL 8 % [mercado2017transthoracic]
We demonstrate the UQ part of the pipeline on a pulmonary hypertensive patient presented in earlier work [pant2022multiscale], whose measured characteristics are presented in Table 6. All measurement errors are assumed independent and uniformly distributed. Sampling was performed using a Sobol‘ sequence, and after 2048 samples parameter input means and variances were deemed converged 12. The achieved samples are then used to solve the forward UQ problem, assuming a shunt placed with the following descriptive parameters of $D_{shunt}\ =\ 1.0\ cm$ and $k_{shunt}\ =\ 3.0\ mL^{2}/(mmHg\cdot s^{2}$). Uncertain predictions are shown in the form of histograms in Figure 13. Note that the predicted outcomes are no longer uniformly distributed, which is due to both the non-linearity of the model input-output relationships, as well as the additional step of solving the inverse problem.
@@ 6.1 Computational time
The computational time of the different metamodels employed is discussed in this section as shown in Table 7. For the building of metamodels, it was observed that neural networks have the fastest training time in the case of a big dataset (>30K>30K data points). The computational time was always in the order of few minutes when training with the full dataset. However, this cost goes to nearly an hour for PCE and around 12 hours for GPs, when running on the same laptop. This is the price paid by GPs in order to offer a natural way to quantify uncertainty in predictions, providing the predicted mean but also a confidence interval, representing the uncertainty at each prediction point. This is particularly valuable in fields like Bayesian optimization, where understanding uncertainty is crucial. It should be noted that when training on a small part of the dataset, 3K points for example, GPs and PCE have training times in the order of minutes.
Once an emulator is trained, obtaining Sobol indices using the Saltelli method [saltelli2004sensitivity] is a relatively fast process for NN and GPs (few minutes), due to the low computational cost per emulator evaluation. PCE has the advantage that Sobol indices are readily available when the model has been trained, with no additional computational cost.
@@
Inverse problem << A minute Requires additional setup Requires additional setup
@@ 6.2 Ensuring physiological validity in metamodels
For prediction purposes, the challenge is to ensure that the outputs of the metamodel adhere to physiological constraints. This can be approached in two main ways. One method involves training the metamodel exclusively on data that is known to fall within physiological ranges. By doing so, the model learns to generate predictions that are inherently physiologically plausible. In our models, we chose to use known ranges of (patho-)physiological behaviour (i.e., pressure, flows and volumes in respective model components) to determine ‘physiological behaviour’, and filter out those samples which are outside of that behaviour [sala2023sensitivity, Haghebaert2025]. However, this approach limits the model’s applicability to the specific ranges of inputs and outputs included in the training data. The second method involves using post-prediction filtering. In this approach, the metamodel generates predictions based on a broader set of data, but these predictions are subsequently checked and corrected to ensure they fall within physiological limits. While this method helps ensure that final outputs are valid, it does not prevent the generation of non-physiological outputs during the modeling process itself.
In the context of SA, which aims to understand how variations in input parameters affect output variability, the approach can differ based on the tools used. One effective strategy proposed in [sala2023sensitivity] is to build a metamodel based on PCE using a training set restricted to input-output pairs that are physiologically valid. This approach ensures that the PCE provides Sobol indices naturally, reflecting the sensitivity of outputs to inputs while avoiding biases from non-physiological data. For those who may not have access to PCE or are unfamiliar with it, another approach is to create a general metamodel, such as one based on GP or NN, using a dataset that includes both physiological and non-physiological pairs. Sobol indices can then be computed using the classical Monte-Carlo approach as shown in Sec. 5.1.
@@ 6.3 Parameter estimation and inverse problems
While outside of the scope of this work, in some cases there are several values or combinations of inputs that lead to the same output, i.e., the chosen set of input parameters is unidentifiable. Depending on the research question, there are several approaches to deal with unidentifiability. First is a Monte Carlo type approach, where the inverse problem is solved several times with different random initializations, leading to a distribution of input parameters [shin2019scalable]. Some methods leverage the sensitivity indices to derive an orthogonal set of input parameters [li2004selection, saxton2023personalised], while other methods allow for a parametric description of non-identifiable parameter combinations [tong2024invaert]. Alternatively, a formal identifiability analysis is performed, iteratively reducing the input space by fixing unidentifiable parameters to produce a reduced, identifiable set; such as profile likelihood analysis [maiwald2016driving, vanlier2012integrated], diaphony [hornfeck2015diaphony, van2020parameter] or determining intra-class correlation [koopsen2024parameter]. However, these methods can be prohibitive due to their computational cost, which can be overcome by using surrogate models.
@@ 6.4 Conclusion
To conclude, overall NN present advantages in terms of implementation and run-time cost-effectiveness to perform the whole pipeline of studying and performing patient-specific modeling.
Code and data availability
Codes and data are available on GitLab at https://gitlab.com/0d-cardiovascular/metamodeling.
Acknowledgments
We acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 864313).
Appendix A Estimation of computational time for full 0D models
In this appendix, we present the estimated computational time for the different tasks using the full 0D models as shown in the table below.
Table 8: Computational cost of different tasks for three models using the original 0D model.
Model/Task 1 Simulation Sensitivity Analysis Inverse Problem Uncertainty Quantification (UQ) UQ of Inverse Problem
Model 1 1.93s in C >>3 days $\sim$20 mins $\sim$8h $\sim$97 days
Model 2 0.3s in Fortran $\sim$4h $\sim$5 mins $\sim$50 mins $\sim$34 days
Model 3 5.6s in Python $\sim$2 days $\sim$5h $\sim$15.5h $\sim$59 weeks